OpenStudy (anonymous):

If the sequence belwo continues according to the pattern shown, what is the sum of first 305 terms of the sequence: -3, 0, 6, -3, 0, 6, -3, 0, 6, -3, 0, 6, ...

OpenStudy (shubhamsrg):

if we divide the series in sets of 3 nos. i,e, {-3,0,6} = 1 set we see each set evolves sum =3 so sum upto first 303 terms i.e. 101 sets = 101 * 3 =303 next 2 terms are -3,0// so final sum is 303-3+0 = 300..which should be the ans..

OpenStudy (anonymous):

Uh, the only answer options given are: a to d viz, -305 305 -610 610

OpenStudy (shubhamsrg):

i'd still stick to my ans..these might be wrong..hmmn..

OpenStudy (anonymous):

the answer is 300

OpenStudy (anonymous):

what @shubhamsrg said is r8